OpenStudy (anonymous):

Find the volume of a cylinder with d = 8.4 m: h = 9.3 m

OpenStudy (anonymous):

55.5896 cubed m. B. 515.123 cubed m C. 67.89 cubed m D. 570.5126 cubed m these are the choices

OpenStudy (anonymous):

\[ V= \pi r^2 h= \pi\frac{ d^2}4h\\ \]

OpenStudy (anonymous):

i dont get it just give me the answer

OpenStudy (akashdeepdeb):

OpenStudy is not a website for answers. Explanation first! And then you yourself can get the answer. :D

OpenStudy (anonymous):

What I gave you is the formula of the volume. You have d, h, so you can compute the volume. You can take \(\pi= 3.14\)